# SONIC: Supersizing Motion Tracking for Natural Humanoid Whole-Body Control

### v1

@@ Abstract
Despite the rise of billion-parameter foundation models trained across thousands of GPUs, similar scaling gains have not been shown for humanoid control. Current neural controllers for humanoids remain modest in size, target a limited behavior set, and are trained on a handful of GPUs over several days. We show that scaling up model capacity, data, and compute yields a generalist humanoid controller capable of creating natural and robust whole-body movements. Specifically, we posit motion tracking as a natural and scalable task for humanoid control, leverageing dense supervision from diverse motion-capture data to acquire human motion priors without manual reward engineering. We build a foundation model for motion tracking by scaling along three axes: network size (from 1.2M to 42M parameters), dataset volume (over 100M frames, 700 hours of high-quality motion data), and compute (9k GPU hours). Beyond demonstrating the benefits of scale, we show the practical utility of our model through two mechanisms: (1) a real-time universal kinematic planner that bridges motion tracking to downstream task execution, enabling natural and interactive control, and (2) a unified token space that supports various motion input interfaces, such as VR teleoperation devices, human videos, and vision-language-action (VLA) models, all using the same policy. Scaling motion tracking exhibits favorable properties: performance improves steadily with increased compute and data diversity, and learned representations generalize to unseen motions, establishing motion tracking at scale as a practical foundation for humanoid control.
@@ 1 Introduction
The past decade has witnessed a remarkable journey in artificial intelligence. The GPT (Achiam et al., 2023) family of models trains on 25,000+ GPUs with trillions of tokens. Video and image-generation models (Ramesh et al., 2022; Rombach et al., 2022; Brooks et al., 2024; Ho et al., 2022; Blattmann et al., 2023) leverage thousands of GPUs processing billions of images. These foundation models have shown a consistent pattern: scale unlocks emergent capabilities, generalization, and robustness that smaller models cannot achieve (Kaplan et al., 2020; Hoffmann et al., 2022; Wei et al., 2022; Bommasani et al., 2021). Yet for sim-to-real humanoid control, similar scaling gains have not been achieved. State-of-the-art humanoid control policies are typically small neural networks – often three-layer MLPs with a few million parameters – trained on a single GPU over a few days for a single task. Even more so, due to the manually engineered reward terms for each task, training a policy for too long may even lead to worse performance (Peng et al., 2018, 2021; Sutton, 2019).
@@
In this work, we address both challenges by identifying motion tracking as the scalable foundation task for humanoid control. Motion tracking leverages human motion capture data, which provides dense, frame-by-frame supervision without reward engineering. Critically, humanoids benefit from decades of motion capture research – datasets covering walking, running, dancing, sports, and object interactions already exist at scale (Mahmood et al., 2019; Punnakkal et al., 2021; Li et al., 2021b). While there exists prior art in motion tracking (Liao et al., 2025; Zhang et al., 2025; Zeng et al., 2025; Yin et al., 2025; Chen et al., 2025; Luo et al., 2023; Wang et al., 2020; He et al., 2025b, 2024), they are mostly limited to showing whole-body motion tracking results on training data and have not demonstrated many downstream tasks beyond motion tracking or navigation. We supersize physics-based motion tracking to unprecedented 100 million frames (at 50 fps) and 128 GPUs training, achieving universal tracking capabilities across diverse human behaviors while maintaining real-time performances. In addition, we show how such a motion tracker can be applied to meaningful downstream tasks, and introduce two key contributions. First, we develop a universal kinematic motion generation system for interactive control, enabling goal-directed tasks such as interactive locomotion and game-like character control through kinematic planning in motion space. Second, we design a universal token space that supports multimodal control – accepting inputs from teleoperation, human videos, music, text, and vision-language-action models through a unified interface (Tevet et al., 2022; Zhang et al., 2023; Kocabas et al., 2020; Li et al., 2021a). This unified framework allows our motion tracker to directly interface with vision-language-action (VLA) models (Bjorck et al., 2025b) and diverse control modalities.
@@
We propose Supersizing mOtion tracking for Natural humanoId Control (SONIC), a framework that enables natural humanoid control across a wide range of applications. We achieve high-precision teleoperation and interactive control capabilities, including running, jumping, and crawling with natural human-like movement. Leveraging our universal token space, our controller can support cross-embodiment motion tracking where estimated whole-body human motion can be directly mapped to humanoid control signals, bypassing the need for retargeting. We demonstrate seamless integration with multi-modal human motion generation models, supporting video, text, and music control through our universal token space. Furthermore, we show that teleoperation data collected through our system can be used to train vision-language-action foundation models, establishing a complete pipeline from motion tracking at scale to foundation model-based humanoid control. These results validate that massive-scale motion tracking serves as a practical and versatile behavioral foundation model for diverse real-world humanoid applications.
@@
  • We identify motion tracking as a scalable foundation task for humanoid control, demonstrating that it exhibits favorable scaling properties with both compute and data diversity and scale up humanoid control to 9k GPU hours and 100 million frames of motion sequences, achieving universal tracking capabilities across diverse human behaviors.
  • We introduce a kinematic motion generation system for interactive control and a universal token space supporting multimodal inputs, including teleoperation, human videos, motion commands, and VLA foundation models, through single-stage training without distillation.
@@ 2.1 Motion Tracking
We employ a comprehensive set of pose-based and physics-based metrics to measure motion imitation performance. The primary measure is the success rate (Succ), where an imitation attempt is deemed unsuccessful if the humanoid deviates too far from the reference motion trajectory. We further report the root-relative mean per-joint position error (MPJPE) $E_{\text{mpjpe}}$ (in mm), quantifying the local accuracy of the imitation. To assess physical fidelity, we also calculate differences in acceleration ($E_{\text{acc}}$, mm/frame2) and velocity ($E_{\text{vel}}$, mm/frame) between the simulated humanoid and the reference human motion.
@@
In Fig.˜2, we analyze the impact of supersizing our humanoid motion tracking system along three key dimensions: GPU hours, model size, and motion dataset size. All evaluations are conducted in Isaac Lab (NVIDIA et al., 2025), with models trained to convergence over 3 to 7 days. For dataset size, we compare with training on LaFAN (0.4M frames), a subset of our in-house dataset (7.4M frames), and our full dataset (100M frames). For GPU hours comparison, we train all models till convergence for training with 8, 32, and 128 GPUs. Across the board, scaling along any of these three axes leads to consistent improvements in motion imitation performance. Notably, increasing the size of the motion dataset yields the most substantial gains, while increased model size and compute (GPU hours) further enhance results. Notice that for GPU hours, parallelizing over more GPUs (e.g., 8 vs 128) is essential, as we notice that training on a smaller number of GPUs yields worse asymptotic performance than a larger number of GPUs. For evaluation, motion imitation is considered unsuccessful if, at any point during tracking, the humanoid’s body height deviates by more than 0.25m from the reference motion or if the root orientation differs by more than 1 radian.
@@
We assess the real-world performance of SONIC by deploying it on 50 diverse motion trajectories, including dance, jumps, and loco-manipulation tasks. As presented in Fig.˜2, our policy achieves motion imitation in the real world that closely matches its simulation results, operating in a true zero-shot manner. Remarkably, it succeeds on all sequences without a single failure (100% success rate), underscoring the robustness and reliability of our tracker in challenging real-world scenarios.
@@ 2.2 Interactive Motion Control
In this section, we showcase the scalability and robustness of SONIC in whole-body, real-time interactive control tasks. We present a kinematic runtime generative motion planner that guides the robot’s policy through user interaction. Our approach employs an autoregressive framework that continually regenerates future kinematic motions conditioned on the previous robot states and incoming user commands. For each planning step, the model generates motion segments lasting between 0.8s and 2.4s, where the duration is automatically determined by the neural planner to maximize flexibility and robustness. The planner achieves inference times under 5 ms on a standard laptop and 12 ms on a Jetson Orin GPU. Replanning is triggered as frequently as every 100 ms, or immediately when user commands are updated, ensuring highly responsive control.
@@
SONIC supports a variety of applications, including but not limited to: (1) navigation control with arbitrary velocity, direction, and style commands; (2) interactive entertainment tasks such as boxing; (3) locomotion skills such as squatting, crawling, kneeling, etc., which are useful for downstream applications like teleoperation. Because both our kinematic planner and tracking policy are trained on the same large-scale dataset. Utilizing the scalable nature of SONIC, we note that all of the applications above were designed afterwards without retraining the planner or the tracking policy. People with minimal animation or programming background can easily design and specify new desired behaviors.
@@
For navigation control, SONIC supports velocity commands ranging from 0.0m/s to 6.0m/s, as well as arbitrary direction commands spanning 0 to 360 degrees. We employ a critically damped spring model to smooth impractical commands that cannot be achieved by the robot within one motion segment. SONIC achieves scalable, responsive and robust navigation control as shown in the first two rows of Fig.˜3. SONIC also supports different styles such as drunken walking, injured walking, happy walking, stealth walking, etc., as shown in the third row of Fig.˜3. The capacity of the model to generate robust inbetweening motions showcases the flexibility of SONIC, and the potential of more natural human-robot interaction.
@@
SONIC further extends its versatility to interactive entertainment tasks such as boxing, as illustrated in the last two rows of Fig.˜3. Existing academic solutions (Starke et al., 2021; Won et al., 2021) and industrial approaches (Unitree, 2025) often utilize a limited collection of boxing clips, requiring switching between multiple expert models or action labels. This approach leads to discontinuous, unnatural transitions and even pauses. SONIC on the other hand enables much more fluid, responsive, and natural motion generation, while retaining the robot’s full freedom of movement throughout the task.
@@
To enable downstream manipulation or navigation in confined environments, skills such as squatting, kneeling, and crawling are essential. As illustrated in Fig.˜4, SONIC supports a rich spectrum of postural modes, including squatting and kneeling, allowing the pelvis height to be smoothly controlled from 0.3m to 0.8m. This greatly enhances the agent’s reachability and adaptability in scenarios such as teleoperation. For navigation in especially tight spaces, SONIC also enables crawling. The robot can move omnidirectionally using its elbows and knees at velocities from 0.0m/s to 0.5m/s.
@@ 2.3 Video Teleoperation and Multi-Modal Cross-Embodiment Control
For music control, the robot generates dance motions conditioned on melodic and rhythmic structure, tracking tempo, and adapting style to musical characteristics.
@@ 2.4.1 VR-Based Whole-Body Teleoperation
We develop a full-body VR teleoperation system using the PICO whole-body motion-tracking interface. This requires wearing the PICO headset, two ankle trackers, and the handheld VR controllers. The PICO interface then provides human motion as full-body human pose estimates in SMPL (Loper et al., 2015) format. The tracked human motion is streamed in real time to our universal control policy, which will be introduced in Sec.˜3.2. The human motion will be encoded via its respective encoder into the universal token space, and then decoded by the robot control decoder to yield low-latency, stable, human-like control, as can be seen in Fig.˜5 (bottom).
@@ 2.4.2 VR-Based 3-Point Teleoperation
To enable scalable and portable data collection for tasks that do not require precise foot placement, we also introduce a lightweight mobile bimanual VR teleoperation interface that operates with the PICO headset and the two handheld controllers (no ankle trackers needed for this mode). The teleoperation interface outputs a compact command consisting of three upper-body SE(3) poses (head and both wrists), finger joint angles, waist height, a locomotion mode (slow walk or fast walk), and a navigation command specifying desired root velocity and heading. These signals are fed into the kinematic motion planner (of Sec.˜3.3) and hybrid encoder (introduced in Sec.˜3.2), then executed by the universal tracking policy.
@@ 2.5 Discussion
Relative to prior trackers (e.g., Any2Track (Zhang et al., 2025), BeyondMimic (Liao et al., 2025), GMT (Chen et al., 2025)), our contributions are threefold: (1) scale: orders-of-magnitude larger training and larger unseen evaluation dataset sets; (2) versatility: robot, human, and hybrid commands in a shared latent, enabling cross-embodiment transfer and versatile interface; and (3) practicality: on-board, real-time sim-to-real control on a Unitree G1 with stronger performance under a common evaluation protocol. Together, these advances move motion tracking from narrow imitation toward a general foundation for whole-body humanoid control.
The planner (Sec.˜2.2) enables responsive navigation, postural mode control (squatting, crawling), and style modulation without retraining. The universal token space (Sec.˜2.3) supports seamless switches among input modalities and cross-embodiment teleoperation. The same interface connects to a vision–language–action model (Sec.˜2.4), providing the system 1 capabilities needed for humanoid loco-manipulation.
Limitations include formal treatment of safety, compliance, and energy efficiency for extended deployments, as well as combating noisy input during deployments. Future work will study scaling laws across more diverse dataset, enable VLA instructed whole-body locomanipulation tasks, and explore joint training of planner, tokenizers, and policy to reduce modality gaps.
In summary, scaling motion tracking yields reliable, general whole-body control; pairing it with a planner, a universal token space, and an efficient onboard stack makes it usable as a system. We expect SONIC to serve as a practical foundation upon which higher-level perception and reasoning can be layered to advance general-purpose humanoid autonomy.

### v2

@@ Abstract
Despite the rise of billion-parameter foundation models trained across thousands of GPUs, similar scaling gains have not been shown for humanoid control. Current neural controllers for humanoids remain modest in size, target a limited set of behaviors, and are trained on a handful of GPUs over several days. We show that scaling up model capacity, data, and compute yields a generalist humanoid controller capable of creating natural and robust whole-body movements. Specifically, we posit motion tracking as a natural and scalable task for humanoid control, leveraging dense supervision from diverse motion-capture data to acquire human motion priors without manual reward engineering. We build a foundation model for motion tracking by scaling along three axes: network size (from 1.2M to 42M parameters), dataset volume (over 100M frames, 700 hours of high-quality motion data), and compute (9k GPU hours). Beyond demonstrating the benefits of scale, we show the practical utility of our model through two mechanisms: (1) a real-time universal kinematic planner that bridges motion tracking to downstream task execution, enabling natural and interactive control, and (2) a unified token space that supports various motion input interfaces, such as VR teleoperation devices, human videos, and vision-language-action (VLA) models, all using the same policy. Scaling motion tracking exhibits favorable properties: performance improves steadily with increased compute and data diversity, and learned representations generalize to unseen motions, establishing motion tracking at scale as a practical foundation for humanoid control.
@@ 1 Introduction
The past decade has witnessed a remarkable journey in artificial intelligence: the GPT (Achiam et al., 2023) family of models are trained on 25,000+ GPUs with trillions of tokens; video and image-generation models (Ramesh et al., 2022; Rombach et al., 2022; Brooks et al., 2024; Ho et al., 2022; Blattmann et al., 2023) leverage thousands of GPUs processing billions of images. These foundation models have shown a consistent pattern: scale unlocks emergent capabilities, generalization, and robustness that smaller models cannot achieve (Kaplan et al., 2020; Hoffmann et al., 2022; Wei et al., 2022; Bommasani et al., 2021). Yet for sim-to-real humanoid control, similar scaling gains have not been achieved. State-of-the-art humanoid control policies are typically small neural networks – often three-layer MLPs with a few million parameters – trained on a single GPU over a few days for a single task. Even more so, due to the manually engineered reward terms for each task, training a policy for too long may even lead to worse performance (Peng et al., 2018, 2021; Sutton, 2019).
@@
In this work, we address both challenges by identifying motion tracking as the scalable foundational task for humanoid control. Motion tracking leverages human motion capture data, which provides dense, frame-by-frame supervision without reward engineering. Critically, humanoids benefit from decades of motion capture research – datasets covering walking, running, dancing, sports, and object interactions already exist at scale (Mahmood et al., 2019; Punnakkal et al., 2021; Li et al., 2021b). While there exists prior art in motion tracking (Liao et al., 2025; Zhang et al., 2025; Zeng et al., 2025; Yin et al., 2025; Chen et al., 2025; Luo et al., 2023; Wang et al., 2020; He et al., 2025b, 2024), they are mostly limited to showing whole-body motion tracking results on training data and have not demonstrated many downstream tasks beyond motion tracking or navigation. We supersize physics-based motion tracking to unprecedented 100 million frames (at 50 fps) and 128 GPUs training, achieving universal tracking capabilities across diverse human behaviors while maintaining real-time performance. In addition, we show how such a motion tracker can be applied to meaningful downstream tasks, and introduce two key contributions. First, we develop a universal kinematic motion generation system for interactive control, enabling goal-directed tasks such as interactive locomotion and game-like character control through kinematic planning in motion space. Second, we design a universal token space that supports multimodal control – accepting inputs from teleoperation, human videos, music, text, and vision-language-action models through a unified interface (Tevet et al., 2022; Zhang et al., 2023; Kocabas et al., 2020; Li et al., 2021a). This unified framework allows our motion tracker to directly interface with vision-language-action (VLA) models (Bjorck et al., 2025b) and diverse control modalities.
@@
We propose Supersizing mOtion tracking for Natural humanoId Control (SONIC), a framework that enables natural humanoid control across a wide range of applications. We achieve high-precision teleoperation and interactive control capabilities, including running, jumping, and crawling with natural human-like movement. Leveraging our universal token space, our controller can support cross-embodiment motion tracking where estimated whole-body human motion can be directly mapped to humanoid control signals, bypassing the need for retargeting. We demonstrate seamless integration with multi-modal human motion generation models, supporting video, text, and music control through our universal token space. Furthermore, we show that teleoperation data collected through our system can be used to train vision-language-action foundation models, establishing a complete pipeline from motion tracking at scale to foundation model-based humanoid control. These results validate that massive-scale motion tracking serves as a practical and versatile foundational task for diverse real-world humanoid applications.
@@
  • We identify motion tracking as a scalable foundational task for humanoid control, demonstrating that it exhibits favorable scaling properties with both compute and data diversity. We scale up humanoid control to 9,000 GPU hours and 100 million frames of motion sequences, achieving universal tracking capabilities across diverse human behaviors.
  • We introduce a kinematic motion generation system for interactive control and a universal token space supporting multimodal inputs, including teleoperation, human videos, motion commands, and VLA foundation models through a single-stage training process without distillation.
@@ 2.1 Motion Tracking
We employ a comprehensive set of pose-based and physics-based metrics to measure motion imitation performance. The primary measure is the success rate (Succ), where an imitation attempt is deemed unsuccessful if the humanoid deviates too far from the reference motion trajectory. We further report the root-relative mean per-joint euclidean position error (MPJPE) $E_{\text{mpjpe}}$ (in mm), quantifying the local accuracy of the imitation. To assess physical fidelity, we also calculate differences in acceleration ($E_{\text{acc}}$, mm/frame2) and velocity ($E_{\text{vel}}$, mm/frame) between the simulated humanoid and the reference human motion.
@@
In Fig.˜2, we analyze the impact of supersizing our humanoid motion tracking system along three key dimensions in simulation and real-world: GPU hours, model size, and motion dataset size. All evaluations are conducted in Isaac Lab (NVIDIA et al., 2025), with models trained to convergence over 3 to 7 days. For dataset size, we compare with training on LaFAN (0.4M frames), a subset of our in-house dataset (7.4M frames), and our full dataset (100M frames). For GPUs hours comparison, we train all models to convergence using 8, 32, and 128 GPUs. Across the board, scaling along any of these three axes leads to consistent improvements in motion imitation performance. Notably, increasing the size of the motion dataset yields the most substantial gains, while increased model size and compute (GPU hours) further enhance results. Notice that, for GPUs hours, parallelizing across more GPUs (e.g., 8 vs 128) is essential, as training on a smaller number of GPUs yields worse asymptotic performance than on a larger number of GPUs. For evaluation, motion imitation is considered unsuccessful if, at any point during tracking, the humanoid’s body height deviates by more than 0.25m from the reference motion or if the root orientation differs by more than 1 radian.
@@
We assess the real-world performance of SONIC by deploying it on 50 diverse motion trajectories, including dance, jumps, and loco-manipulation motions. As presented in Fig.˜2, our policy achieves motion imitation in the real world that closely matches its simulation results, operating in a true zero-shot manner. Remarkably, it succeeds on all sequences without a single failure (100% success rate), underscoring the robustness and reliability of our tracker in challenging real-world scenarios.
@@ 2.2 Interactive Motion Control
In this section, we showcase the scalability and robustness of SONIC in whole-body, real-time interactive control tasks. We present a kinematic runtime generative motion planner that guides the robot’s policy through user interaction. Our approach employs an autoregressive framework that continually regenerates future kinematic motions conditioned on the previous states and incoming user commands. For each planning step, the model generates motion segments lasting between 0.8s and 2.4s, where the duration is automatically determined by the neural planner to maximize flexibility and robustness. The planner achieves inference times under 5 ms on a standard laptop and 12 ms on a Jetson Orin GPU. Replanning is triggered as frequently as every 100 ms, or immediately when user commands are updated, ensuring highly responsive control.
@@
SONIC supports a variety of applications, including but not limited to: (1) navigation control with arbitrary velocity, direction, and style commands; (2) interactive entertainment tasks such as boxing; (3) locomotion skills such as squatting, crawling, kneeling, etc., which are useful for downstream applications like teleoperation. Because both our kinematic planner and tracking policy are trained on the same large-scale dataset. Utilizing the scalable nature of SONIC, we note that all the applications above were designed afterwards without retraining the planner or the tracking policy. People with minimal animation or programming backgrounds can easily design and specify new desired behaviors.
@@
For navigation control, SONIC supports velocity commands ranging from 0.0m/s to 6.0m/s, as well as arbitrary direction commands spanning 0 to 360 degrees. We employ a critically damped spring model to smooth impractical commands that cannot be achieved by the robot within one motion segment. SONIC achieves scalable, responsive, and robust navigation control, as shown in the first two rows of Fig.˜3. SONIC also supports different styles such as drunken walking, injured walking, happy walking, stealth walking, etc., as shown in the third row of Fig.˜3. The capacity of the model to generate robust inbetweening motions showcases the flexibility of SONIC, and the potential of more natural human-robot interaction.
@@
SONIC further extends its versatility to interactive entertainment tasks such as boxing, as illustrated in the last two rows of Fig.˜3. Existing academic solutions (Starke et al., 2021; Won et al., 2021) and industrial approaches (Unitree, 2025) often utilize a limited collection of boxing clips, requiring a switch between multiple expert models or action labels. This approach leads to discontinuous, unnatural transitions and even pauses. SONIC, on the other hand, enables much more fluid, responsive, and natural motion generation while retaining the robot’s full freedom of movement throughout the task.
@@
To enable downstream manipulation or navigation in confined environments, skills such as squatting, kneeling, and crawling are essential. As illustrated in Fig.˜4, SONIC supports including squatting, kneeling, and crawling. For squatting and kneeling, we slow the pelvis height to be smoothly controlled from 0.3m to 0.8m. For navigation in especially tight spaces, SONIC also enables crawling. The robot can move omnidirectionally using its elbows and knees at velocities from 0.0m/s to 0.5m/s.
@@ 2.3 Video Teleoperation and Multi-Modal Cross-Embodiment Control
For music control, GENMO generates dance motions (which our tracking policy imitates) conditioned on melodic and rhythmic structure, tracking tempo, and adapting style to musical characteristics.
@@ 2.4.1 VR-Based Whole-Body Teleoperation
We develop a full-body VR teleoperation system using the PICO whole-body motion-tracking interface Zhao et al. (2025). This requires wearing the PICO headset, two ankle trackers, and the handheld VR controllers. The PICO interface then provides human motion as full-body human pose estimates in SMPL (Loper et al., 2015) format. The tracked human motion is streamed in real time to our universal control policy, which will be introduced in Sec.˜3.2. The human motion will be encoded via its respective encoder into the universal token space, and then decoded by the robot control decoder to yield low-latency, stable, human-like control, as can be seen in Fig.˜5 (bottom).
@@ 2.4.2 VR-Based 3-Point Teleoperation
To enable scalable and portable data collection for tasks that do not require precise foot placement, we also introduce a lightweight mobile bimanual VR teleoperation interface that operates with the PICO headset and two handheld controllers (no ankle trackers are needed for this mode). The teleoperation interface outputs a compact command consisting of three upper-body SE(3) poses (head and both wrists), finger joint angles, waist height, a locomotion mode (slow walk or fast walk), and a navigation command specifying the desired root velocity and heading. These signals are fed into the kinematic motion planner (of Sec.˜3.3) and hybrid encoder (introduced in Sec.˜3.2), and then executed by the universal tracking policy.
@@ 2.5 Discussion
Relative to prior trackers such as Any2Track (Zhang et al., 2025), BeyondMimic (Liao et al., 2025), and GMT (Chen et al., 2025), our system advances the state of the art along three dimensions: (i) we train on motion data that is orders of magnitude larger than existing systems and evaluate on substantially broader, previously unseen datasets; (ii) via a shared latent command space that unifies hybrid input commands to enable multi-modal control using a unified tracking policy; and (iii) our universal kinematic planner that bridges the gap between real-world user-demands and our tracking policies. These components collectively move motion tracking from demonstration-specific imitation toward a general foundation for humanoid control.
Limitations include the formal treatment of safety, compliance, and energy efficiency for extended deployments on the humanoid, as well as handling noisy input during deployments. Future work will study scaling laws across more diverse dataset, enable VLA instructed whole-body locomanipulation tasks, and explore joint training of planner, tokenizer, and policy to reduce modality gaps.
In summary, scaling motion tracking yields reliable, general whole-body control; pairing it with a planner, a universal token space, and an efficient onboard stack makes it usable as a system. We expect SONIC to serve as a practical foundation upon which higher-level perception and reasoning can be built to advance general-purpose humanoid autonomy.